# ResoFilter: Fine-grained Synthetic Data Filtering for Large Language Models through Data-Parameter Resonance Analysis

### v1

@@ 4.1 Experiments Setup
Model Method 25% 50% 75% Full SFT Time Cost(Hour)
Gemma2-2B Random 0.6042 0.6254 0.6474 0.649 -
@@
Table 1: Figure 1 presents a comprehensive comparison of our proposed method against several baseline approaches across three different model architectures: Gemma2-2B, Llama2-7B, and Llama2-13B. In addition to the three most commonly used types of methods, we also cited two methods for data selection Li et al. (2024c) Li et al. (2024b). As can be seen, our performance has achieved the best among these methods with only a small increase in time cost compared to other techniques.
@@ 4.2 Main Results
Fields Method 25% 50% 75% Full SFT
CodeX Random 0.2979 0.3549 0.404 0.4098
Loss 0.3146 0.3442 0.3872
Our method 0.3579 0.425 0.3966
MMLU Random 0.5105 0.5152 0.4858 0.4658
Loss 0.507 0.501 0.4961
Our method 0.5129 0.4893 0.4952
BBH Random 0.3796 0.3759 0.3824 0.3704
Loss 0.3796 0.3667 0.3685
Our method 0.3861 0.362 0.3778
Table 3: Cross-domain evaluation results comparing our proposed method with random sampling. There has been significant improvement in code generation while in general domain it is not obvious. We have more detail experiments and explanations in Appendix B
@@ B.2 Performance on BBH
Appendix C Additional Experiments on Model Scalability and Reproducibility
To demonstrate the robustness and scalability of our method, we conducted two sets of additional experiments: (1) evaluating performance on larger models, and (2) verifying reproducibility through multiple runs.
C.1 Experiments on Larger Models
We validated our method’s effectiveness on larger and more recent models, including Gemma2-9B and Llama2-70B. Importantly, we used data filtered by smaller models (Gemma2-2B and Llama2-7B respectively) to train these larger models, demonstrating the practical efficiency of our "filter-once-apply-many" approach.
Model Base Score Method 25% 50% 75%
Gemma2-9B 0.69 Random 0.7498 0.7467 0.7483
Ours 0.7596 0.7566 0.7741
Llama2-70B 0.61 Random 0.7671 0.7839 0.7960
Ours 0.7733 0.7877 0.8059
These results demonstrate that our method maintains its effectiveness when scaled to larger models, with consistent improvements over random selection across different model sizes.
C.2 Reproducibility Analysis

### v3

@@ 3.3 ResoFilter
Module Selection
Based on our comprehensive analysis (detailed in Appendix D), we identified $W_{up}$ as the most effective weight module for data filtering. As shown in Table 8, $W_{up}$ consistently outperforms other modules across different data ratios, with particularly strong performance at 50% and 75% data volumes. This selection is further supported by ablation studies demonstrating the module’s stability and effectiveness in capturing semantic changes during fine-tuning.
@@ 4.1 Experiments Setup
Category Model/Field Method 25% 50% 75% Full SFT Time*
Models Gemma2-2B Random 0.6042 0.6254 0.6474 0.649 -
@@
Domains CodeX Random 0.2979 0.3549 0.404 0.4098
Loss 0.3146 0.3442 0.3872
Our method 0.3579 0.425 0.3966
MMLU Random 0.5105 0.5152 0.4858 0.4658
Loss 0.507 0.501 0.4961
Our method 0.5129 0.4893 0.4952
BBH Random 0.3796 0.3759 0.3824 0.3704
Loss 0.3796 0.3667 0.3685
Our method 0.3861 0.362 0.3778
Table 1: Combined results of cross-model and cross-domain evaluations. The upper section compares ResoFilter with baselines across model architectures (Gemma2-2B, Llama2-7B, Llama2-13B), while the lower section demonstrates domain generalization (CodeX, MMLU, BBH). Time* indicates training time cost (hours) per method, which is only applicable to model architecture experiments.
@@ 4.2 Main Results
Scalability to Larger Models
Model Base Score Method 25% 50% 75%
Gemma2-9B 0.69 Random 0.7498 0.7467 0.7483
Ours 0.7596 0.7566 0.7741
Llama2-70B 0.61 Random 0.7671 0.7839 0.7960
Ours 0.7733 0.7877 0.8059
To specifically validate our method’s scalability, we conducted extended experiments on Gemma2-9B and Llama2-70B. Using the identical data filtered by their smaller counterparts (Gemma2-2B and Llama2-7B respectively), these larger models achieve comparable relative improvements to smaller models (3.2% vs. 3.4% average gain at p25). This demonstrates that: 1) The data selection criteria learned from smaller models generalize effectively to larger architectures; 2) Our method scales naturally with model capacity without requiring additional filtering iterations. The consistent performance across orders-of-magnitude parameter differences (2B to 70B) suggests our approach captures fundamental data quality characteristics rather than model-specific artifacts.
@@ B.2 Performance on BBH
Appendix C Reproducibility Analysis